Dear all,
please find below a short summary of the activities of the th related to the AO system.


People at LBTO:


R. Briguglio


reote support
A. Riccardi and M.Xompero (VNC+Skype)


Summary
Telescope operations began at 23:36UT. During the first part of the night, approximately until 4:30+1UT, several problems slew down the activity:
* act#455 and #255 suffered from miscalibration and were therefore excluded;
*act #363, #366, #465 were exluded from ICC actuators list
*software error (wrong configuration file), corrected
*the "ear" on the shell edge moved from a maximum gap of 60um to a minimum of 39um (nominal rest position)
*shell ripped by the internal control loop, several times.

The shell was successfully set at 4:30+1UT. During operations, a single restart was needed to re-align the finite state machine status, that were mismatching from the true system status because an operator's error. Since 4:30UT until the end of operations (the dome was closed and the shell put on its rest position at 13:00UT) no fault occurred. High order modes offload and astigmatism automatic correction were working all the night on.

A deep analysis of the initial faults is on progress and will be discussed in the next reports. System temperatures were recorded on a spreadsheet file:
20100331 Time (UT) Elevation Flux Chamber Inlet Outlet Hub Coldplate Refplate Outside

4.59 61.00 9.89 3.95 3.66 5.51 4.10 4.13 4.69 3.70

5.21 62.00 9.87 3.81 3.61 5.67 3.99 3.91 4.31 3.60

6.10 79.00 9.88 3.35 3.90 5.78 3.50 4.00 4.00 3.50

6.40 60.00 9.89 2.98 3.96 5.95 3.37 4.02 3.96 2.20

7.30 66.00 9.90 2.44 3.77 5.84 2.60 4.06 3.71 2.80

8.55 49.00 9.91 1.29 3.85 6.05 1.67 3.99 3.52 0.50

9.31 54.00 9.88 0.79 3.96 6.11 1.51 4.11 3.47 0.50

11.02 48.00 9.89 0.87 3.85 5.97 1.02 3.98 3.42 0.10

11.31 78.00 9.81 0.69 3.58 5.56 1.05 3.64 3.22 0.05

12.16 47.00 9.84 0.86 3.00 5.24 0.62 3.29 2.83 -0.20

12.51 90.00 9.81 0.30 2.90 5.18 0.92 3.16 2.79 -0.30

A more detailed report may be found here: TelLogLBT672a20100330


Ciao

Runa

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r2 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback